OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-5, -1) and is perpendicular to y = 5 Over 2x + 2.

OpenStudy (anonymous):

Perpendicular line given is : \[y = \frac{5}{2}x + 2\] Right ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you tell me what is the slope of this line ??

OpenStudy (anonymous):

Just compare it with: \[y = mx + c\]

OpenStudy (anonymous):

srry but i dont understand how to do it completly

OpenStudy (anonymous):

Just tell the slope.. Compare it with the equation that I have written above. Simply tell me what is the coefficient of x there ??

OpenStudy (anonymous):

srry i just dont know could u work it out with give me an example???

OpenStudy (anonymous):

by giving me an example?

OpenStudy (anonymous):

Ok I give you one example: Suppose you have: \[y = \frac{6}{7}x + 2\] Compare this equation with: \[y = mx + c\] They are just the same equations; On comparing you get m as ; \(\frac{6}{7}\) and c as 2. Now can you do the same for : \[y = \frac{5}{2}x + 2\] Can you tell what is m here and c here ??

OpenStudy (anonymous):

5 over 2x is M and 2 is c

OpenStudy (anonymous):

Here m is only 5/2..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This is called slope..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now one thing you must remember that: In case of perpendicular lines, if one has slope m then the other line will have the slope of \(\frac{-1}{m}\)..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, here you got the slope has 5/2. So the line perpendicular to it will have slope of \(\large \frac{-2}{5}\).. Getting ??

OpenStudy (anonymous):

sorta

OpenStudy (anonymous):

I show you..

OpenStudy (anonymous):

the only part i dont understand is the -1/m part

OpenStudy (anonymous):

\[\large \frac{-1}{m} = \frac{-1}{\frac{5}{2}} \implies \frac{-2}{5}\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so ur the flipping the 5 and 2 and makeing the 2 negative right or is there more to it?

OpenStudy (anonymous):

In case of perpendicular lines, their slope product is -1.. \[m_1m_2 = -1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

From here: \(m_1 = \frac{5}{2}\) So, you will get \(m_2 = \frac{-2}{5}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So now you have slope and one point is also given to you: The required equation is given by: \[\large y - y_1 = m(x - x_1)\] Here: \[x_1 = -5, y_1 = -\] \[m = \frac{-2}{5}\] Just plug in the values and find the equation..

OpenStudy (anonymous):

\[y_1 = -1\]

OpenStudy (anonymous):

im lost now :PP

OpenStudy (anonymous):

I give you one example: Suppose the points given will be: \((-4, -5)\) And the slope is \(\frac{-7}{6}\) So according to that equation: \[y - y_1 = m(x - x_1)\] \[y - (-5) = \frac{-7}{6}(x - (-4))\] \[y + 5 = \frac{-7}{6}(x +4)\] \[6y + 30 = -7x - 28\] Do the same above..

OpenStudy (anonymous):

would it be y = -2/5x-3?